OpenStudy (anonymous):

please help me. algebra 2 test on wensday!!!

OpenStudy (anonymous):

\[(5x+1)^{1/4}+4\]

OpenStudy (anonymous):

i meant to put equals 4

OpenStudy (loser66):

\[\large (5x+1)^{\frac{1}{4}} = 4\] right?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

exponent to the 4 both sides \[\large ((5x+1)^{\frac{1}{4}})^4 = 4^4\] you get 5x +1 = 256 now it is easy to solve for x, right?

OpenStudy (anonymous):

yes but how did u get 256 for 4 to the 4th power

OpenStudy (loser66):

4*4*4*4 = 256

OpenStudy (anonymous):

ohhhhhh thank u lol

OpenStudy (loser66):

ok